OpenStudy (anonymous):

Which is a ordered solution to the pair? 7x + 2y= -39 -8x +3y= -3 please tutor me

OpenStudy (anonymous):

\[ Which ordered pair is a solution of the system? 7x + 2y = –39 –8x + 3y = –3\]

OpenStudy (anonymous):

7x+2y=-39 Simple and best practice solution for 7x+2y=-39 equation. Check how easy it is, and learn it for the future. Our solution is simple, and easy to understand, so don`t hesitate to use it as a solution of your homework. If it's not what You are looking for type in the equation solver your own equation and let us solve it. Equation: Solution for 7x+2y=-39 equation: Simplifying 7x + 2y = -39 Solving 7x + 2y = -39 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-2y' to each side of the equation. 7x + 2y + -2y = -39 + -2y Combine like terms: 2y + -2y = 0 7x + 0 = -39 + -2y 7x = -39 + -2y Divide each side by '7'. x = -5.571428571 + -0.2857142857y Simplifying x = -5.571428571 + -0.2857142857y

OpenStudy (anonymous):

Solution for -8x+3y=-3 equation: Simplifying -8x + 3y = -3 Solving -8x + 3y = -3 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-3y' to each side of the equation. -8x + 3y + -3y = -3 + -3y Combine like terms: 3y + -3y = 0 -8x + 0 = -3 + -3y -8x = -3 + -3y Divide each side by '-8'. x = 0.375 + 0.375y Simplifying x = 0.375 + 0.375y